APPLICATIONS OF SPFIT TO DIATOMIC ROVIBRONIC SPECTRA

\underline{BRIAN J. DROUIN}, CHARLES E. MILLER, HERBERT M. PICKETT, and EDWARD A. COHEN, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr, Pasadena, CA 91109-8099.

Applications of SPFIT^{a} to fitting the spectra of diatomic molecules will be demonstrated. Examples will include: fitting of multiple isotopomers and vibronic levels with mass-independent Dunham parameters including Born-Oppenheimer corrections; fitting transitions between states of different multiplicity; simultaneous fitting of data from microwave, infrared and optical experiments weighted by appropriate experimental uncertainties; and simultaneous fitting of data sets containing both resolved and unresolved hyperfine structure.

aH. M. Pickett, J. Mol. Spectrosc. 148, 371-377 (1991)